# Mixture-of-Recursions: Learning Dynamic Recursive Depths for Adaptive Token-Level Computation

### v1

@@ 1 Introduction
One proven route to parameter efficiency is layer tying, in which a shared set of weights is reused across multiple layers (Dehghani et al., 2018; Lan et al., 2019; Gholami and Omar, 2023; Bae et al., 2024; Takase and Kiyono, 2021). For adaptive computation, a common approach is early-exiting, which dynamically allocates compute by exiting earlier in the network when predicting simpler tokens (Elhoushi et al., 2024; Schuster et al., 2022; Elbayad et al., 2020; Bae et al., 2023). Despite the progress achieved along each of these individual efficiency axes, an architecture that effectively unifies both parameter efficiency and adaptive computation is still missing. Recursive Transformers (Bae et al., 2024; Fan et al., 2024; Giannou et al., 2023; Yang et al., 2023; Saunshi et al., 2025; Geiping et al., 2025), models that repeatedly apply the same set of shared layers multiple times, offer a strong foundation due to their built-in weight sharing. However, prior attempts at dynamic recursion have often been constrained by practical hurdles, such as requiring additional specialized training procedures or facing challenges in efficient deployment. This has led most approaches to still default to a simpler fixed-depth recursion, which applies the same amount of computation to every token and is thus incapable of delivering truly adaptive token-level compute allocation.
@@ 2.1 Preliminary
We examine four parameter-sharing strategies: Cycle, Sequence, and their variants Middle-Cycle and Middle-Sequence. Table 1 summarizes two main designs, and the full list is provided in Table 5 in the Appendix. In Cycle sharing, recursion blocks are reused cyclically. For example, consider an original non-recursive model with L=9L{=}9 layers and its recursive counterpart using $N_{r}{=}3$ recursions. Under the “Cycle” strategy, the layers are shared and unrolled as [$(0,1,2),(0,1,2),(0,1,2)$]. In “Sequence” sharing, each recursion block reuses the same layer consecutively before moving to the next, resulting in [$(0,0,0),(1,1,1),(2,2,2)$] for the same configuration. Both have the same effective number of layers when unrolled (L=9L{=}9), but with a different order. Furthermore, the “Middle” variants preserve full-capacity parameters at the first and last layers ($\Phi_{0}$ and $\Phi_{L-1}$), while sharing weights among the intermediate layers.
@@ 2.2 Mixture-of-Recursions
To ensure coherent progression through steps, we adopt hierarchical filtering: only tokens selected at recursion step $r$ can be re-evaluated at r+1r{+}1. This simulates early-exit behavior while learning from scratch. As deeper layers tend to encode increasingly abstract and sparse information (Li et al., 2022; Yang et al., 2024; Nawrot et al., 2024), this mechanism prioritizes computation for only the most demanding tokens.
@@
[…]
@@ 3 Experiments
[…]
@@
$\text{Recursive}^{\dagger}$ - - M-Cyc 2 167M 16.5 20B 2.8079 31.0 37.1 66.7 52.3 40.8 27.5 42.6
@@ 3.2 IsoFLOP Analysis
As shown in Figure 3, MoR consistently outperforms recursive baselines across all model sizes and compute budgets. While it underperforms the vanilla model at the smallest model size (135M)—likely due to a recursive capacity bottleneck—this gap closes rapidly at scale. For >360M parameters, MoR not only matches but often exceeds the Vanilla Transformer, particularly under low and mid-range budgets. Overall, these results highlight that MoR is a scalable and efficient alternative to standard Transformers. It achieves strong validation performance with significantly lower parameter counts, making it a strong candidate for both pretraining and large‑scale deployment. Further details are presented in Appendix C.
@@ 3.3 Inference Throughput Evaluation
[…]
@@ 5.1 Compute-optimal Scaling Analysis
As illustrated in 5(a), MoR exhibits a distinct compute-optimal scaling behavior compared to baselines under isoFLOPs constraints. The larger slope (i.e., closer to zero) of MoR’s optimal path indicates that it benefits more significantly from increases in parameter count (i.e., less data-hungry). This is likely because the performance of the shared parameter block itself becomes important, even more than feeding in additional data. Therefore, the optimal scaling policy for MoR models favors allocating resources to increasing model capacity by using larger models trained for shorter steps.
@@ 6 Related Work
Parameter sharing provides an orthogonal path to efficiency (Dehghani et al., 2018; Lan et al., 2019; Xia et al., 2019; Takase and Kiyono, 2021; Bae et al., 2024; Jaegle et al., 2021). The Universal Transformer first showed that repeatedly applying a single block can match the representational power of deep, non-shared stacks (Dehghani et al., 2018). Looped Transformers shown to be effective, can act as programmable computers (Giannou et al., 2023), learn iterative data-fitting algorithms (Yang et al., 2023), generalize to much longer inputs on algorithmic tasks (Fan et al., 2024), and illuminate few-shot learning by mimicking multi-step optimizers (Gatmiry et al., 2024). Furthermore, Bae et al. (2024) mitigate the accuracy loss often associated with weight tying by adding low-rank adaptation (LoRA) adapters (Hu et al., 2022) in each loop, yielding Relaxed Recursive Transformers. Recent work further demonstrates that Recursive Transformers excel at latent reasoning via recurrent depth (Geiping et al., 2025). While most prior studies focus on the efficiency gains from weight tying, the recursive architecture itself offers a second level: inspired by early-exiting (Schuster et al., 2022) and compute routing (Raposo et al., 2024), one can vary the number of recursions per input (e.g., per token), allocating compute only where it is most beneficial during both training and inference.
@@
Many works have shown that dynamic compute allocation can markedly reduce the cost of training and inference, from traditional neural networks (Bengio et al., 2015; Huang et al., 2016; Teerapittayanon et al., 2016; Panda et al., 2016) to large language models (Hou et al., 2020; Elbayad et al., 2020; Fedus et al., 2022; Bae et al., 2023; Elhoushi et al., 2024; Raposo et al., 2024). Early exiting methods learn to halt processing for “easy” samples (e.g., tokens or sequences in language modeling) by skipping the remaining layers (Elbayad et al., 2020; Schuster et al., 2022; Dehghani et al., 2018; Mofakhami et al., 2024). Alternatively, early exits can be combined with speculative decoding techniques (Chen et al., 2023; Leviathan et al., 2023) during inference by leveraging lower layers for fast drafting (Bae et al., 2023; Elhoushi et al., 2024). Recently, Mixture-of-Depths (MoD) (Raposo et al., 2024) reframed adaptive depth as a routing problem: a router at each layer selects a subset of tokens to receive the full computation, while the rest bypass the layer, yielding finer-grained conditional compute. This new form of adaptive allocation is well suited to Transformer architectures and has already been extended to other modalities (Zhang et al., 2024a; Luo et al., 2024), highlighting a promising paradigm of dynamic compute at token-level granularity. MoR applies this routing idea to recursive Transformers: tokens are dynamically sent through repeated calls of a single, weight‑tied block instead of through distinct layers. This shift keeps parameter count constant, allows arbitrarily deep (adaptive) compute beyond the model’s physical depth.
@@ 7.1 Limitations and Future Works
Expansion to multimodal and non‑text domains.
@@ 7.2 Acknowledgements
We thank Jacob Eisenstein for valuable feedback on an earlier version of the paper. We also thank the Google Cloud Platform for awarding Google Cloud credits for this project.
@@ Appendix C Expanded Results of IsoFLOP Analysis
In the main paper (§3.2), we compared Vanilla, Recursive and our Mixture‑of‑Recursions (MoR) models under matched training compute. Four base model capacities were studied—135M, 360M, 730M and 1.7B parameters. For recursive and MoR models, we fix the recursion count to $N_{r}\!=\!3$, so the number of unique parameters is roughly one‑third of the vanilla counterpart. Each architecture is trained once for the largest compute budget (16.5EB)6661EB = $10^{18}$ floating‑point operations. and the resulting checkpoint is re‑used to obtain the 5EB and 2EB variants, as detailed below.
@@
Trapezoid learning‑rate schedule with checkpoint reuse.
@@
$\eta(t)=\begin{cases}\frac{t}{w}\,\eta_{\max},&0\leq t<w\quad\text{(warm‑up)},\\[4.0pt] \eta_{\max},&w\leq t<p\quad\text{(plateau)},\\[2.0pt] \eta_{\max}\!\Bigl(1-\tfrac{t-p}{d}\Bigr),&p\leq t<p+d\quad\text{(cool‑down)},\end{cases}$
where $w$ denotes the warm‑up interval, pwp-w is the constant‑LR plateau, and $d$ represents the cool-down segment. This stable phase allows us to efficiently manage experiments by saving intermediate checkpoints and then running additional cool-down steps from those points, according to each budget. For the warm‑up we allocate $5\,\%$ of the total training steps of the smallest budget (2EB), and we set the cool‑down steps to $20\,\%$ of the total training steps for each corresponding budget.
@@
Table 7 reports NLL on FineWeb-Edu validation set and few‑shot accuracy on six benchmarks. Our findings reveal a clear trend: more compute consistently leads to better models, evidenced by lower NLL and improved accuracy with higher FLOPs. However, weight-sharing alone in recursive models degraded performance compared to Vanilla, a clear trade-off for the reduced parameters. Crucially, token-routed MoR models overcome this shortcoming, catching up to and then surpassing Vanilla models from 360M parameters upward, all while utilizing only one-third of the parameters. This performance advantage persists at 730M and 1.7B parameter scales.
Table 7: Detailed results of isoFLOP analysis across three compute budgets. We evaluate negative log‑likelihood (NLL) on the FineWeb‑Edu validation set and few‑shot accuracy on six downstream tasks for four base model sizes (135M, 360M, 730M, 1.7B). Each model was initially trained up to 16.5EB and sliced back to 5EB and 2EB via mid‑training checkpoints using a trapezoid learning‑rate schedule. All models used three recursion steps. For MoR models, we use expert-choice routing and recursion-wise caching mechanisms. We highlight the best-performing model in each setting in gray.
@@ Appendix E Expanded Results of Parameter Sharing Strategy
This section complements the ablation in §4.1 by providing the full quantitative panorama behind 4(b). We revisit the four weight‑tying schemes—Cycle, Sequence, Middle‑Cycle, and Middle‑Sequence—on two base model scales (135M and 360M non‑embedding parameters) and two different recursion depths ($N_{r}=2$ and 3). All models were trained from scratch for 10B tokens under identical optimization hyperparameters. Validation NLL on FineWeb‑Edu and averaged few‑shot accuracy over six benchmarks are summarized in Table 8.
Middle‑Cycle is consistently the safest choice.
@@
Figure 6: Validation negative log‑likelihood (lower is better) on FineWeb‑Edu for four parameter‑sharing strategies. Bars are grouped by model capacity (135M vs. 360M parameters). Middle‑Cycle consistently attains the lowest NLL, with its margin widening as either model size or depth increases. The horizontal dashed lines mark the untied (non-sharing) baselines: the lower red line represents the full capacity model, while the upper black line represents a parameter-matched reduced model with a footprint equal to the unique trainable parameter sizes of the recursive model.
Behavior under continued pre‑training (up‑training).
Table 9 extends the study by “up‑training” models—continuing from open-sourced SmolLM (Allal et al., 2024) checkpoints for an additional 5B tokens. Both Middle strategies demonstrate superior performance across all settings, and notably, they significantly outperform the reduced baseline models that are initialized in the same manner but without recursion. The other strategies reach a performance plateau earlier, suggesting that they have limited room for further improvement in capacity.
@@ F.2 Router Performance Evaluation Metrics
[…]
@@ G.2 Performance Comparison of KV Sharing Strategy
[…]

### v2

@@ 1 Introduction
One proven route to parameter efficiency is layer tying, in which a shared set of weights is reused across multiple layers (Dehghani et al., 2018; Lan et al., 2019; Takase and Kiyono, 2021; Gholami and Omar, 2023; Bae et al., 2024). For adaptive computation, a common approach is early-exiting, which dynamically allocates compute by exiting earlier in the network when predicting simpler tokens (Elhoushi et al., 2024; Schuster et al., 2022; Elbayad et al., 2020; Bae et al., 2023). Despite the progress achieved along each of these individual efficiency axes, an architecture that effectively unifies both parameter efficiency and adaptive computation is still missing. Recursive Transformers (Bae et al., 2024; Fan et al., 2024; Giannou et al., 2023; Yang et al., 2023; Saunshi et al., 2025; Geiping et al., 2025; Aleksandrov et al., 2025), models that repeatedly apply the same set of shared layers multiple times, offer a strong foundation due to their built-in weight sharing. However, prior attempts at dynamic recursion have often been constrained by practical hurdles, such as requiring additional specialized training procedures or facing challenges in efficient deployment. This has led most approaches to still default to a simpler fixed-depth recursion, which applies the same amount of computation to every token and is thus incapable of delivering truly adaptive token-level compute allocation.
@@ 2.1 Preliminary
We examine four parameter-sharing strategies: Cycle, Sequence, and their variants Middle-Cycle and Middle-Sequence. Table 1 summarizes two main designs, and the full list is provided in Table 5 in the Appendix. In Cycle sharing, recursion blocks are reused cyclically. For example, consider an original non-recursive model with $L{=}9$ layers and its recursive counterpart using $N_{r}{=}3$ recursions. Under the “Cycle” strategy, the layers are shared and unrolled as [$(0,1,2),(0,1,2),(0,1,2)$]. In “Sequence” sharing, each recursion block reuses the same layer consecutively before moving to the next, resulting in [$(0,0,0),(1,1,1),(2,2,2)$] for the same configuration. Both have the same effective number of layers when unrolled ($L{=}9$), but with a different order. Furthermore, the “Middle” variants preserve full-capacity parameters at the first and last layers ($\Phi_{0}$ and $\Phi_{L-1}$), while sharing weights among the intermediate layers.
@@ 2.2 Mixture-of-Recursions
To ensure coherent progression through steps, we adopt hierarchical filtering: only tokens selected at recursion step $r$ can be re-evaluated at $r{+}1$. This simulates early-exit behavior while learning from scratch. As deeper layers tend to encode increasingly abstract and sparse information (Li et al., 2022; Yang et al., 2024; Nawrot et al., 2024), this mechanism prioritizes computation for only the most demanding tokens.
@@
$\mathcal{H}_{t}^{r+1}=\begin{cases}g_{t}^{r}f(\mathcal{H}_{t}^{r},\,\Phi^{\prime})+\mathcal{H}_{t}^{1},&\text{if }r=i\\ g_{t}^{r}f(\mathcal{H}_{t}^{r},\,\Phi^{\prime}),&\text{otherwise}\\ \end{cases}$ (2.2)
@@ 3 Experiments
Table 3: Comparison of MoR, Recursive, and Vanilla Transformers under both fixed FLOPs (16.5e18) and token (20B) settings. All models are trained on FineWeb-Edu and evaluated by validation negative log-likelihood (NLL) and few-shot accuracy. For the isoFLOP rows, the number of training tokens ($N_{tok}$) varies by model efficiency. For the fixed-token rows, we report the effective FLOPs consumed. For the model sizes, we report non-embedding parameter counts. For the KV mechanisms, we distinguish between Cache (recursion-wise caching) and Share (recursive sharing). $\text{}^{\dagger}$n recursive models, all tokens go through fixed recursion depths ($N_{r}$), instead of adaptive depths.
@@
$\text{Recursive}^{\dagger}$ M-Cyc 2 167M 16.5 20B 2.8079 31.0 37.1 66.7 52.3 40.8 27.5 42.6
@@ 3.2 IsoFLOP Analysis
As shown in Figure 3, MoR consistently outperforms recursive baselines across all model sizes and compute budgets. While it underperforms the vanilla model at the smallest model size (135M)—likely due to a recursive capacity bottleneck—this gap closes rapidly at scale. For >360M parameters, MoR not only matches but often exceeds the Vanilla Transformer, particularly under low and mid-range budgets. Overall, these results highlight that MoR is a scalable and efficient alternative to standard Transformers. It achieves strong validation performance with significantly lower parameter counts, making it a strong candidate for both pretraining and large-scale deployment. Further details are presented in Appendix C.
@@ 3.3 Inference Throughput Evaluation
In 4(a), across both batch settings, all MoR variants outperform the vanilla baseline, which even leverages continuous sequence-wise batching (Yu et al., 2022; Kwon et al., 2023). Increasing recursion depth leads to more tokens exiting early and a further reduction in KV cache usage. This, in turn, boosts throughput significantly (e.g., MoR-4 achieves up to a 2.06$\times$ speedup with $B=\text{Max}$). While there’s a slight performance degradation, it can be a favorable trade-off given the substantial throughput gain. These results support that the integration of the depth-wise batching paradigm with early-exiting can significantly accelerate MoR’s actual deployment throughput.
@@ 5.1 Compute-optimal Scaling Analysis
As illustrated in 5(a), MoR exhibits a distinct compute-optimal scaling behavior compared to baselines under isoFLOPs constraints. The flatter slope of MoR’s optimal path (a line connecting stars) indicates that it benefits more significantly from increases in parameter count (i.e., less data-hungry). This is likely because the performance of the shared parameter block itself becomes important, even more than feeding in additional data. Therefore, the optimal scaling policy for MoR models favors allocating resources to increasing model capacity by using larger models trained for shorter steps.
@@ 6 Related Work
Parameter sharing provides an orthogonal path to efficiency (Dehghani et al., 2018; Lan et al., 2019; Xia et al., 2019; Jaegle et al., 2021; Takase and Kiyono, 2021; Ng and Wang, 2024; Bae et al., 2024). The Universal Transformer first showed that repeatedly applying a single block can match the representational power of deep, non-shared stacks (Dehghani et al., 2018). Looped Transformers shown to be effective, can act as programmable computers (Giannou et al., 2023), learn iterative data-fitting algorithms (Yang et al., 2023), generalize to much longer inputs on algorithmic tasks (Fan et al., 2024), and illuminate few-shot learning by mimicking multi-step optimizers (Gatmiry et al., 2024). Furthermore, Bae et al. (2024) mitigate the accuracy loss often associated with weight tying by adding low-rank adaptation (LoRA) adapters (Hu et al., 2022) in each loop, yielding Relaxed Recursive Transformers. Recent work further demonstrates that Recursive Transformers excel at latent reasoning via recurrent depth (Geiping et al., 2025). While most prior studies focus on the efficiency gains from weight tying, the recursive architecture itself offers a second level: inspired by early-exiting (Schuster et al., 2022) and compute routing (Raposo et al., 2024), one can vary the number of recursions per input (e.g., per token), allocating compute only where it is most beneficial during both training and inference.
@@
Many works have shown that dynamic compute allocation can markedly reduce the cost of training and inference, from traditional neural networks (Bengio et al., 2015; Huang et al., 2016; Teerapittayanon et al., 2016; Panda et al., 2016) to large language models (Hou et al., 2020; Elbayad et al., 2020; Fedus et al., 2022; Bae et al., 2023; Elhoushi et al., 2024; Raposo et al., 2024). Early exiting methods learn to halt processing for “easy” samples (e.g., tokens or sequences in language modeling) by skipping the remaining layers (Elbayad et al., 2020; Schuster et al., 2022; Dehghani et al., 2018; Mofakhami et al., 2024). Alternatively, early exits can be combined with speculative decoding techniques (Chen et al., 2023; Leviathan et al., 2023) during inference by leveraging lower layers for fast drafting (Bae et al., 2023; Elhoushi et al., 2024). Recently, Mixture-of-Depths (MoD) (Raposo et al., 2024) reframed adaptive depth as a routing problem: a router at each layer selects a subset of tokens to receive the full computation, while the rest bypass the layer, yielding finer-grained conditional compute. This new form of adaptive allocation is well suited to Transformer architectures and has already been extended to other modalities (Zhang et al., 2024a; Luo et al., 2024), highlighting a promising paradigm of dynamic compute at token-level granularity. MoR applies this routing idea to recursive Transformers: tokens are dynamically sent through repeated calls of a single, weight-tied block instead of through distinct layers. This shift keeps parameter count constant, allows arbitrarily deep (adaptive) compute beyond the model’s physical depth.
@@ 7.1 Limitations and Future Works
Expansion to multimodal and non-text domains.
@@ 7.2 Acknowledgements
We thank Jacob Eisenstein for valuable feedback on an earlier version of the paper. We also thank Seungyeon Kim, Mostafa Elhoushi, Sangdoo Yun for helpful conversations. Finally, we thank the Google Cloud Platform for awarding Google Cloud credits for this project.
@@ Appendix C Expanded Results of IsoFLOP Analysis
In the main paper (§3.2), we compared Vanilla, Recursive and our Mixture-of-Recursions (MoR) models under matched training compute. Four base model capacities were studied—135M, 360M, 730M and 1.7B parameters. For recursive and MoR models, we fix the recursion count to $N_{r}\!=\!3$, so the number of unique parameters is roughly one-third of the vanilla counterpart. Each architecture is trained once for the largest compute budget (16.5EB)6661EB = $\,{=}\,10^{18}$ floating-point operations. and the resulting checkpoint is re-used to obtain the 5EB and 2EB variants, as detailed below.
@@
Trapezoid learning-rate schedule with checkpoint reuse.
@@
$\eta(t)=\begin{cases}\frac{t}{w}\,\eta_{\max},&0\leq t<w\quad\text{(warm\mbox{-}up)},\\[4.0pt] \eta_{\max},&w\leq t<p\quad\text{(plateau)},\\[2.0pt] \eta_{\max}\!\Bigl{(}1-\tfrac{t-p}{d}\Bigr{)},&p\leq t<p+d\quad\text{(cool\mbox{-}down)},\end{cases}$
where $w$ denotes the warm-up interval, $p-w$ is the constant-LR plateau, and $d$ represents the cool-down segment. This stable phase allows us to efficiently manage experiments by saving intermediate checkpoints and then running additional cool-down steps from those points, according to each budget. For the warm-up we allocate $5\,\%$ of the total training steps of the smallest budget (2EB), and we set the cool-down steps to $20\,\%$ of the total training steps for each corresponding budget.
@@
Table 7 reports NLL on FineWeb-Edu validation set and few-shot accuracy on six benchmarks. Our findings reveal a clear trend: more compute consistently leads to better models, evidenced by lower NLL and improved accuracy with higher FLOPs. However, weight-sharing alone in recursive models degraded performance compared to Vanilla, a clear trade-off for the reduced parameters. Crucially, token-routed MoR models overcome this shortcoming, catching up to and then surpassing Vanilla models from 360M parameters upward, all while utilizing only one-third of the parameters. This performance advantage persists at 730M and 1.7B parameter scales.
Table 7: Detailed results of isoFLOP analysis across three compute budgets. We evaluate negative log-likelihood (NLL) on the FineWeb-Edu validation set and few-shot accuracy on six downstream tasks for four base model sizes (135M, 360M, 730M, 1.7B). Each model was initially trained up to 16.5EB and sliced back to 5EB and 2EB via mid-training checkpoints using a trapezoid learning-rate schedule. All models used three recursion steps. For MoR models, we use expert-choice routing and recursion-wise caching mechanisms. We highlight the best-performing model in each setting in gray.
@@ Appendix E Expanded Results of Parameter Sharing Strategy
This section complements the ablation in §4.1 by providing the full quantitative panorama behind 4(b). We revisit the four weight-tying schemes—Cycle, Sequence, Middle-Cycle, and Middle-Sequence—on two base model scales (135M and 360M non-embedding parameters) and two different recursion depths ($N_{r}=2$ and 3). All models were trained from scratch for 10B tokens under identical optimization hyperparameters. Validation NLL on FineWeb-Edu and averaged few-shot accuracy over six benchmarks are summarized in Table 8.
Middle-Cycle is consistently the safest choice.
@@
Figure 6: Validation negative log-likelihood (lower is better) on FineWeb-Edu for four parameter-sharing strategies. Bars are grouped by model capacity (135M vs. 360M parameters). Middle-Cycle consistently attains the lowest NLL, with its margin widening as either model size or depth increases. The horizontal dashed lines mark the untied (non-sharing) baselines: the lower red line represents the full capacity model, while the upper black line represents a parameter-matched reduced model with a footprint equal to the unique trainable parameter sizes of the recursive model.
Behavior under continued pre-training (up-training).
Table 9 extends the study by “up-training” models—continuing from open-sourced SmolLM (Allal et al., 2024) checkpoints for an additional 5B tokens. Both Middle strategies demonstrate superior performance across all settings, and notably, they significantly outperform the reduced baseline models that are initialized in the same manner but without recursion. The other strategies reach a performance plateau earlier, suggesting that they have limited room for further improvement in capacity.
@@ F.2 Router Performance Evaluation Metrics
$\mathrm{MaxVio}=\frac{\max_{i}\mathrm{Load}_{i}-\overline{\mathrm{Load}}_{i}}{\overline{\mathrm{Load}}_{i}},$
@@ G.2 Performance Comparison of KV Sharing Strategy
Table 12: Comparison of KV cache sharing strategies across Vanilla, Recursive, and MoR Transformers. Models are pretrained on 10B tokens of FineWeb-Edu, and evaluated using negative log-likelihood (NLL) on train set and few-shot accuracy across benchmarks. KV sharing denotes use of recursive KV sharing mechanism. If MoR is mentioned without further specification of KV sharing strategy, it implies the use of the recursion-wise caching strategy. $\text{}^{\dagger}$t indicates training with hybrid KV sharing that leverages shared caches for inactive positions while updating active ones through actual computation.